OpenStudy (anonymous):

what the equation OF THIS LINE (4,5) (-2,-3)

OpenStudy (anonymous):

y-y0=m(x-x0) first we find m=deltay/deltX ==>M=8/6 THEN place into the equation: y-5=8/6(x-4)

OpenStudy (midhun.madhu1987):

 2 point form ofma line